### v1

@@ 3.2 Assumption smuggling
Empirical studies often contain a section on “mechanisms” with empirical tests that purport to elucidate why a causal effect exists without laying out the assumptions these tests require. This is a prime example of assumption smuggling: attempting to maintain nominally weak assumptions with empirical tests that actually require a stronger set of unmentioned assumptions. Of course, like with unattended luggage in an airport, most instances of assumption smuggling are unintentional. Many strong assumptions have been embedded in the parametric models commonly used in political science, and many of us have built our statistical intuitions from that foundation. But one benefit of the recent “credibility revolution” has been to encourage researchers to be more clear about the assumptions of their research designs. Our goal is to help mediation-skeptic scholars who wish to test causal mechanisms do so with a clear articulation of what exactly they are assuming.
@@ 4 Intermediate Outcome Tests
Scholars using IOT usually present comparisons of $\tau$ and $\alpha$ as a “necessary but not sufficient argument” for causal mechanisms. Suppose, for example, we have estimated a positive ATE, $\widehat{\tau}>0$. We might believe that a certain mediator captures the mechanism of interest in the sense that a positive indirect effect through it explains the positive $\widehat{\tau}$. If indirect effects are effects along the path $A_{i}\rightarrow M_{i}\rightarrow Y_{i}$, then we might believe that a positive average treatment effect on the mediator would establish the first of these arrows. Under this argument, a positive $\widehat{\alpha}$ is necessary but not sufficient since we also require information about the $M_{i}\rightarrow Y_{i}$ arrow. Green, Ha and Bullock (2010, p. 207) explain the logic of this approach (see Callis, Dunning and Tuñón (2024, pp. 61–2) for a similar description):
@@
In other words, under this argument, we might view IOTs as a kind of falsification test—we hope they have the potential to rule out mechanisms such that passing the test has some informative value. Scholars often take this strategy explicitly because they are uncomfortable with the assumptions required to estimate the causal relationship between the mediator and the outcome.
@@
Unfortunately, the “necessary but not sufficient” argument above has been shown to be invalid. Glynn (2012) showed that the product of the average effect of $A_{i}$ on $M_{i}$ and the average effect of $M_{i}$ on $Y_{i}$ does not produce estimates of indirect effects when treatment effects can vary from unit to unit. [^3] Indeed, Green, Ha and Bullock (2010) recognize this when they say, “this kind of analysis makes some important assumptions about homogeneous treatment effects, but the point is that this type of exploratory investigation may provide some useful clues to guide further experimental investigation.” We have not found researchers using IOTs to either acknowledge these assumptions or be as cautious in the interpretation of IOTs as Green, Ha and Bullock (2010) recommend. We extend those results by providing nonparametric bounds on what we can learn about the indirect effect from the maintained assumptions of the mediation skeptic.
@@ 4.1 Principal stratification approach to mediation
We borrow the terminology of compliers and defiers from the literature on instrumental variables to mean units that comply with or defy the proposed causal mechanism (rather than treatment assignment as in the instrumental variables setting). We can define an individual-level indirect effect as
@@
$\alpha$ $=\mathbb{E}[M_{i}(1)-M_{i}(0)]=1\cdot\rho_{10}+(-1)\cdot\rho_{01}+0\cdot\rho_{11}+0\cdot\rho{00}=\rho_{10}-\rho_{01},$
where the second equality follows from an application of the iterated expectations. The always-takers and never-takers contribute nothing because their values of $M_{i}$ never change with treatment.
@@
$=\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)|M_{i}(1)=1,M_{i}(0)=0]\cdot\rho_{10}$
$\qquad-\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)|M_{i}(1)=0,M_{i}(0)=1]\cdot\rho_{01}.$ (4)
The last expression shows that the ANIE is a function of the average effect of the mediator on the outcome for the mechanism compliers and defiers. The compliers contribute positive effects of the mediator while defiers contribute their negative effects.
The IOT approach restricts itself from identifying the ATM, or $\rho_{10}-\rho_{01}$ in the principal strata framework. With no further assumptions, what can that tell us about mechanisms? From (4), we can see that we can have a non-zero ANIE even with an ATM of exactly zero. In particular, the ATM could be zero if there are equal numbers of mechanism compliers and defiers—that is, if an equal number of respondents’ mediators are positively and negatively affected by treatment—but the effect of $M_{i}$ on $Y_{i}$ is different for these groups. Thus, IOTs alone cannot rule out an indirect effect without further assumptions, leaving their interpretation as falsification tests suspect. Furthermore, non-zero ATMs are also not sufficient for indirect effects because any difference between $\rho_{10}$ and $\rho_{01}$ in (4) can be offset by differences in the effect of $M_{i}$ on $Y_{i}$ in each of these groups. Thus, when we maintain the set of mediation-skeptical assumptions, intermediate outcomes provide almost no information about causal mechanisms, at least in the form of indirect effects.
@@ 4.2 The hidden assumptions of intermediate outcome tests
In words, Assumption 3 states that an individual’s potential mediator value is always weakly greater if they were to receive the treatment. This rules out the existence of mediator-defiers in the population (that is, units for which $M_{i}(1)=0,M_{i}(0)=1$), meaning that $\rho_{01}=0$ and the ATM will be equal to $\rho_{10}$. Applying this assumption to (4) allows us to write the ANIE as
@@
i.e., a scaled (by $\rho_{10}$) controlled direct effect (CDE) of $M_{i}$ on $Y_{i}$ among the mechanism compliers. Thus, under MMR, we recover a weak “product rule” interpretation of the indirect effect. We can write the ANIE as the product of the effect of the treatment on the mediator and the effect of the mediator on the outcome among units whose mediator is affected by treatment, or
@@
Thus, under MMR, an average effect of treatment on the moderator of zero rules out an average indirect effect, and the IOT can now properly function as a falsification test. Unfortunately, intermediate outcome tests remain insufficient for establishing causal mediation because even if the ATM is non-zero, the effect of $M_{i}$ on $Y_{i}$ for the mechanism compliers may still be zero. Is the MMR assumption plausible? The answer to this question will, of course, depend on the context. There may be many settings where this is a perfectly reasonable assumption. A survey experiment with a prompt designed to anger respondents is unlikely to make some respondents less angry. The difficulty with this assumption is that it must be true for all units to make the above interpretations correct, which may be difficult to sustain in many cases.
@@ 4.3 Sharp bounds for the indirect effect
The previous section showed when IOTs can be used to rule out causal mechanisms. These results, however, are somewhat broad, and it would be useful to have a more precise sense of what we can learn about the indirect effect from a given empirical setting. We next turn to establishing nonparametric bounds for the indirect effect, which are the values of an indirect effect that are consistent with a set of empirical results and assumptions (Manski, 1995). To do so, we use the linear programming approach of Balke and Pearl (1997) and Sachs et al. (2023). [^4] We ignore covariates in these derivations for presentational simplicity, but it is possible to incorporate covariates for either identification or for narrowing the bounds (Levis et al., 2023).
Let $p_{ym\cdot a}=\mathbb{P}(Y_{i}=y,M_{i}=m\mid A_{i}=a)$ be the joint distribution of the outcome and the mediator within levels of treatment. Imai, Tingley and Yamamoto (2013) derived the sharp nonparametric bounds for the ANIE under the mediation-skeptic assumptions (that is, Assumption 1) as
These bounds tell us the most we can learn about the ANIE from the joint distribution of the data (including the mediator) under the mediation-skeptical assumptions. Note that because $p_{ym\cdot a}$ is nonnegative, zero will always be contained in these bounds, which is consistent with the earlier discussion that IOTs are uninformative about indirect effects under the mediation-skeptical assumptions. With these bounds, we cannot rule out a mechanism even if the ATM is zero since non-zero indirect effects will always be in the bounds, no matter the values of the ATM. Thus, we are extremely limited in what we can learn about indirect effects from IOTs.
We can derive sharp bounds for the indirect effect under different assumptions. The bounds under MMR provide bounds that we can write in terms of the ATM.
@@
What assumptions would be necessary to allow an IOT to establish the presence of an indirect effect? Any such assumption would need to place restrictions on the causal relationship between the mediator and the outcome. Indeed, as shown by Pearl (2001) and Imai, Keele and Yamamoto (2010), we must assume that the mediator is as-if randomized with respect to the outcome (conditional on pretreatment covariates) to point identify the ANIE. Of course, the mediation-skeptical approach seeks to avoid any such assumptions.
@@
Proposition 2 shows that even if we assume monotonic mediator response and a positive average treatment effect of the mediator on the outcome, we still cannot learn the direction of the indirect effect. Equation 6 shows why this is the case—the indirect effect focuses on the effect of $M_{i}$ on $Y_{i}$ for the mechanism compliers, whereas the average treatment effect of $M_{i}$ on $Y_{i}$ is the effect averaged across all units. Thus, even under these assumptions, IOTs cannot be sufficient tests for causal mechanisms.

### v2

@@ 3.2 Assumption smuggling
Empirical studies often contain a section on “mechanisms” with empirical tests that purport to elucidate why a causal effect exists without laying out the assumptions these tests require. These mechanism sections are often a prime example of assumption smuggling: attempting to maintain nominally weak assumptions with empirical tests that actually require a stronger set of unmentioned assumptions. Of course, like with unattended luggage in an airport, most instances of assumption smuggling are unintentional. Parametric models commonly used in political science often embed many strong assumptions, and many of us have built our statistical intuitions from that foundation. But one benefit of the recent “credibility revolution” has been to encourage researchers to be more explicit about the assumptions of their research designs. Our goal is to help mediation-skeptic scholars who wish to test causal mechanisms do so with a clear articulation of what exactly they are assuming.
@@ 4.1 What are IOTs and how are they interpreted?
IOTs then proceed to interpret estimates of this ATM in terms of how it relates to causal mechanisms, generally, or average indirect effect specifically. There are two interpretations of the IOT results that differ in their strength:
  1. 1.
    Falsification interpretation: no ATM ($\alpha=0$) implies $M_{i}$ is likely not part of an indirect effect/causal mechanism.
  2. 2.
    Establishment interpretation: a non-zero ATM ($\alpha\neq 0$) is suggestive evidence that $M_{i}$ is part an indirect effect/causal mechanism.
To be clear, the logic of IOTs is often left implicit in papers that deploy them, but their interpretation usually falls into one (or both) of these two buckets.
The falsification interpretation views IOTs as “necessary but not sufficient evidence” for causal mechanisms. In our empirical application on prejudice reduction, the authors find that the perspective-getting intervention has a negative overall treatment effect on prejudicial attitudes, or $\widehat{\tau}$ < 0. We might suspect that a measure of empathy for the outgroup could be part of the causal mechanism here in the sense that perspective-getting increases empathy, which in turn decreases prejudice. Taken together, these effects would imply a negative indirect effect that explains the ATE. The falsification interpretation says that if we do not see an average effect of the perspective getting on empathy, then empathy is unlikely to explain the causal effect because any indirect effects through empathy must go along the path $A_{i}\rightarrow M_{i}\rightarrow Y_{i}$. Green, Ha and Bullock (2010, p. 207) explain the logic of the falsification interpretation (see Callis, Dunning and Tuñón (2024, pp. 61–2) for a similar description):
@@
In other words, under this argument, researchers view IOTs as a kind of falsification test—they hope they have the potential to rule out mechanisms such that passing the test has some informative value. Scholars often take this strategy explicitly because they are uncomfortable with the assumptions required to estimate the causal relationship between the mediator and the outcome.
The establishment interpretation goes even further and interprets nonzero ATMs as suggestive evidence for a causal mechanism. Often, scholars test mediators that they assume or suspect have a particular causal effect on the outcome. For example, in our second empirical application on transitional justice and democratic attitudes, the authors interpret the significant effect of visiting the transitional justice museum as suggestive evidence of mechanism, operating under the assumption that emotions like fear of repressive regimes lead to a higher level of support for democracy (Balcells, Palanza and Voytas, 2022, p.506).
4.2 The pitfalls of interpreting IOTs
What do we actually learn when we perform IOTs? With a binary outcome, we know that an indirect effect must logically be between $-1$ and $1$ before we see any data. We can use the technique of nonparametric bounds to determine what values of an indirect effect are consistent with a set of empirical results and a set of maintained assumptions (Manski, 1995). To do so, we use the linear programming approach of Balke and Pearl (1997) and Sachs et al. (2023). [^3] We ignore covariates in these derivations for presentational simplicity, but it is possible to incorporate covariates for either identification or for narrowing the bounds (Levis et al., 2023).
Let $p_{ym\cdot a}=\mathbb{P}(Y_{i}=y,M_{i}=m\mid A_{i}=a)$ be the joint distribution of the outcome and the mediator within levels of treatment. In the context of our running example, this would be the joint distribution of prejudice and reactive empathy within levels of the perspective-getting treatment. Imai, Tingley and Yamamoto (2013) derived the sharp nonparametric bounds for the ANIE under the mediation-skeptic assumptions (that is, Assumption 1) as
We refer to these as the randomization bounds. These bounds tell us the most we can learn about the ANIE from the joint distribution of the data (including the mediator) under the mediation-skeptical assumptions. Because the ATM is a deterministic function of the joint distribution, these bounds also tell us all that we can learn about the ANIE from an IOT.
What can we learn from these sharp randomization bounds? Unfortunately, because $p_{ym\cdot a}$ is nonnegative, the randomization bounds will always contain positive values, negative values, and a precisely zero ANIE. This result implies that, under randomization alone, IOTs cannot support the falsification or establishment interpretations. In short, IOTs under randomization alone do not allow us to learn about the presence or sign of the ANIE. However, we do learn something from the data. Logically, we know that the indirect effect must be between -1 and 1 for a binary outcome. The randomization bounds improve over those simple logical constraints and will rule out extreme values of the ANIE in either direction.
Why does randomization alone fail to justify the key IOT interpretations? There are two broad problems that IOTs face under the mediation-skeptical assumptions. First, we often lack credible causal evidence for the relationship between the mediator and the outcome, undermining the basis of the establishment interpretation. In the substantive example, scholars are apprehensive about interpreting any relationship between reactive empathy and prejudicial views as the causal effect of the former on the latter, even conditional on covariates. While it may be possible to rely on previous studies or experimental manipulation of $M_{i}$ in a separate arm to learn its average causal effect, we show below that knowledge of the average $M_{i}\rightarrow Y_{i}$ effect does not provide much information about the causal mechanisms.
@@ 4.2 The pitfalls of interpreting IOTs
The second and more fundamental problem with IOTs is that treatment heterogeneity between units renders average effects relatively uninformative for mediation quantities. For example, Glynn (2012) showed that the product of the average effect of $A_{i}$ on $M_{i}$ and the average effect of $M_{i}$ on $Y_{i}$ does not produce valid estimates of indirect effects when treatment effects can vary from unit to unit, undermining the idea that a zero ATM implies zero ANIE. [^4] Indeed, Green, Ha and Bullock (2010) recognize this when they say, “this kind of analysis makes some important assumptions about homogeneous treatment effects, but the point is that this type of exploratory investigation may provide some useful clues to guide further experimental investigation.” We have not found researchers using IOTs to either acknowledge these assumptions or be as cautious in the interpretation of IOTs as Green, Ha and Bullock (2010) recommend. Returning to the prejudice application, we may estimate no effect of the perspective-getting intervention on the reactive empathy mediator, but this may mask positive and negative effects for equal-sized groups in the population, which average out to an overall null effect. In this setting, many values of indirect effect are possible because it may be the case that the group for whom perspective getting increases empathy has a positive effect of empathy on prejudice, which would lead to a positive indirect effect rather than the theorized negative indirect effect.
Treatment effect heterogeneity is pernicious because these types of unknown correlations undermine both the falsification and establishment interpretation. Detecting a significant nonzero average treatment effect on the mediator is uninformative because we learn nothing about the correlation between the individual-level effects of $A_{i}$ on $M_{i}$ and the individual-level effects of $M_{i}$ on $Y_{i}$. Below, we develop a formal framework to show how these two problems manifest and derive bounds that show precisely what we learn about indirect effects under various assumptions.
@@ 4.3 Principal stratification approach to mediation
We borrow the terminology of compliers and defiers from the literature on instrumental variables to mean units that comply with or defy the proposed causal mechanism (rather than treatment assignment as in the instrumental variables setting). In the prejudice-reduction application, mechanism compliers would be the respondents for whom the perspective-getting intervention increases reactive empathy, whereas mechanism defiers are those for whom the intervention decreases reactive empathy.
We can define an individual-level indirect effect as
@@
$\alpha$ $=\mathbb{E}[M_{i}(1)-M_{i}(0)]=1\cdot\rho_{10}+(-1)\cdot\rho_{01}+0\cdot\rho_{11}+0\cdot\rho_{00}=\rho_{10}-\rho_{01},$
where the second equality follows from an application of the iterated expectations. In other words, the ATM in the prejudice-reduction experiment would be the proportion of respondents whose reactive empathy increases because of treatment minus the proportion of respondents whose reactive empathy decreases because of treatment. The always-takers and never-takers contribute nothing because their values of $M_{i}$ never change with treatment.
@@
$=\underbrace{\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)|M_{i}(1)=1,M_{i}(0)=0]}_{M\to Y\text{ effect for mechanism compliers}}\cdot\rho_{10}$
$\qquad-\underbrace{\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)|M_{i}(1)=0,M_{i}(0)=1]}_{M\to Y\text{ effect for mechanism defiers}}\cdot\rho_{01}.$ (4)
The last expression shows that the ANIE is a function of the average effect of the mediator on the outcome for the mechanism compliers and defiers. The compliers contribute positive effects of the mediator while defiers contribute their negative effects. In the prejudice-reduction context, these are the causal effects of reactive empathy on prejudicial attitudes for those whose reactive empathy is positively (compliers) or negatively (defiers) affected by treatment.
The IOT approach restricts itself from identifying the ATM, or $\rho_{10}-\rho_{01}$ in the principal strata framework. With no further assumptions, what can that tell us about mechanisms? From (4), we can see that we can have a nonzero ANIE even with an ATM of exactly zero. In particular, the ATM could be zero if there are equal numbers of mechanism compliers and defiers—that is, if an equal number of respondents’ mediators are positively and negatively affected by treatment—but the effect of $M_{i}$ on $Y_{i}$ is different for these groups. Thus, IOTs alone cannot rule out an indirect effect without further assumptions, leaving their interpretation as falsification tests suspect. Furthermore, nonzero ATMs are also not sufficient for indirect effects because any difference between $\rho_{10}$ and $\rho_{01}$ in (4) can be offset by differences in the effect of $M_{i}$ on $Y_{i}$ in each of these groups. Thus, when we maintain the set of mediation-skeptical assumptions, intermediate outcomes provide relatively little information about causal mechanisms, at least in the form of indirect effects.
The principal strata formulation of the ANIE in (4) can also clarify how the mediation assumptions identify the ANIE. One consequence of the mediator as-if randomization in Assumption 2 is that the effects of treatment on the mediator and the effect of the mediator on the outcome are statistically independent. Thus, under this strong assumption, knowing that one is a mechanism complier or defier does not change the average effect of reactive empathy on prejudice. Omitting covariates for clarity, we can then simplify (4) as
$\delta(a)$ $=\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)|M_{i}(1)=1,M_{i}(0)=0]\cdot\rho_{10}$
$\qquad-\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)|M_{i}(1)=0,M_{i}(0)=1]\cdot\rho_{01}.$
$=\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)]\cdot\rho_{10}-\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)]\cdot\rho_{01}$
$=\underbrace{\mathbb{E}[Y_{i}(a,1)-Y_{i}(a,0)]}_{M\to Y\text{ effect}}\cdot\underbrace{(\rho_{10}-\rho_{01})}_{A\to Y\text{ effect}}.$
Thus, the mediation assumptions imply powerful restrictions on effect heterogeneity, allowing us to use the product rule to identify the ANIE. Of course, the strength of this assumption is also why so many are skeptical of the assumption.
@@ 4.4 The hidden assumptions of intermediate outcome tests
In words, Assumption 3 states that an individual’s potential mediator value is always weakly greater if they were to receive the treatment. This assumption rules out the existence of mediator-defiers in the population (that is, units for which $M_{i}(1)=0,M_{i}(0)=1$), meaning that $\rho_{01}=0$ and the ATM will be equal to $\rho_{10}$. In the prejudice-reduction context, MMR requires that there be no respondent whose reactive empathy is lower because of treatment. Kwon and Roth (2024) leverage this type of monotonicity assumption to develop a test of the sharp null that $M_{i}$ explains all of the treatment effect, $Y_{i}(a,m)=Y_{i}(m)$ for all $i$, $a$, and $m$. While this is an important test, we believe there are many causal effects that have many mechanisms (some of which are unmeasured), making the sharp null a less interesting test in these contexts.
Applying this assumption to (4) allows us to write the ANIE as
@@
i.e., a scaled (by $\rho_{10}$) controlled direct effect (CDE) of $M_{i}$ on $Y_{i}$ among the mechanism compliers. Thus, we recover a weak “product rule” interpretation of the indirect effect under MMR. We can write the ANIE as the product of the effect of the treatment on the mediator and the effect of the mediator on the outcome among units whose mediator is affected by treatment, or
@@
Thus, under MMR, an average effect of treatment on the moderator of zero rules out an average indirect effect, and the IOT can now properly function as a falsification test. Unfortunately, intermediate outcome tests remain insufficient for establishing causal mediation because even if the ATM is nonzero, the effect of $M_{i}$ on $Y_{i}$ for the mechanism compliers may still be zero.
Is the MMR assumption plausible? The answer to this question will, of course, depend on the context. There may be many settings where this is a perfectly reasonable assumption. A survey experiment with a prompt designed to anger respondents is unlikely to make some respondents less angry. The difficulty with this assumption is that it must be true for all units to make the above interpretations correct, which may be difficult to sustain in many cases. Below, we discuss one way to test the assumption.
@@ 4.5 Sharp bounds for the indirect effect
The previous section showed how to use IOTs to rule out causal mechanisms when the effect of the treatment on the mediator is precisely zero. What about situations where the ATM is small but nonzero? We now derive sharp bounds for the indirect effect under different assumptions. The first result imposes MMR and provides bounds in terms of the ATM.
@@
Proposition 2 shows that even if we assume monotonic mediator response and a positive average treatment effect of the mediator on the outcome, we still cannot learn the direction of the indirect effect. Equation (6) shows why this is the case—the indirect effect focuses on the effect of $M_{i}$ on $Y_{i}$ for the mechanism compliers, whereas the average treatment effect of $M_{i}$ on $Y_{i}$ is the effect averaged across all units. Thus, IOTs cannot be sufficient tests for causal mechanisms even under these assumptions.
Lastly, is there a scenario under which IOTs do constitute sufficient evidence for causal mechanisms? In Supplemental Materials C, we find that a full monotonicity assumption—namely, that all causal effects are in a single direction at the individual level—can allow us to establish the ANIE when the IOTs are nonzero. Of course, to make these restrictions, we would have to know the effect of the mediator on the outcome, which is often unrealistic.
Given the importance of the MMR assumption, how can we determine whether it holds? Testing MMR is challenging because it restricts individual-level treatment effects, which are unidentified. We can, however, exploit what MMR implies for other quantities. In particular, MMR states that individual’s treatment effect of $A_{i}$ on $M_{i}$ ($M_{i}(1)-M_{i}(0)$) be greater than zero, which implies that any conditional average treatment effects (CATEs) of the treatment on the mediator must be positive or equal to zero (i.e., $\mathbb{E}[M_{i}(1)-M_{i}(0)|X_{i}=x]\geq 0$). CATEs capture the average treatment effect of $A_{i}$ on $M_{i}$ for different subgroups of the data, defined by covariates $X_{i}$. Fortunately, the CATEs, unlike the individual-level treatment effects, are identified under some restrictions, allowing us to conduct a falsification test. We can use modern machine learning methods such as causal forests (Wager and Athey, 2018) to estimate the distribution of CATEs for a given set of covariates. If any of the estimates are below zero, this indicates a violation of the MMR assumption. Of course, even if all CATE estimates are above zero, this does not rule out all possible violations of MMR: there may still be individuals within each subgroup with negative individual effects that are canceled out by those with positive effects.
What if we cannot sustain the MMR assumption but wish to impose some structure beyond randomization on our problem? In Supplemental Materials B, we derive the sharp bounds for the ANIE under an assumed upper bound on the proportion of mechanism defiers. These bounds would allow researchers to conduct a sensitivity analysis to see how their inferences depend on this quantity.